# Demystifying Quantum Power Flow: Unveiling the Limits of Practical Quantum Advantage

### v2

@@ Abstract
Quantum computers hold promise for solving problems intractable for classical computers, especially those with high time and/or space complexity. The reduction of the power flow (PF) problem into a linear system of equations, allows formulation of quantum power flow algorithms, based on quantum linear system solving methods such as the Harrow-Hassidim-Lloyd (HHL) algorithm. The speedup due to QPF algorithms is often claimed to be exponential when compared to classical PF solved by state-of-the-art algorithms. We investigate the potential for practical quantum advantage in solving QPF compared to classical methods on gate-based quantum computers. We meticulously scrutinize the end-to-end complexity of QPF approach, providing a nuanced evaluation of the purported quantum speedup in this problem. Our analysis establishes a best-case bound for the HHL-based quantum power flow complexity, conclusively demonstrating that HHL-based method has higher runtime complexity compared to classical algorithm for solving the direct current power flow (DCPF) and fast decoupled load flow problem. Additionally, we establish that for potential PQA to exist it is necessary to consider DCPF-type problems with a very narrow range of condition number values and readout requirements.
The authors acknowledge the funding provided by LANL’s Directed Research and Development (LDRD) project: “High-Performance Artificial Intelligence”(20230771DI) and The Department of Energy (DOE), USA under the Advanced Grid Modeling (AGM) program. The research work conducted at Los Alamos National Laboratory is done under the auspices of the National Nuclear Security Administration of the U.S. Department of Energy under Contract No. 89233218CNA000001
@@ I Introduction
The simultaneous task of providing electrical energy to the world and fighting climate change by switching transportation to electricity will require electricity generation to increase by more than two and half times by 2050, from 2021 levels [1]. This increased reliance on electricity will make secure power grid operation an even more important challenge. Modern power system operation heavily relies on the ability of system operators to compute the system state accurately and efficiently. The fundamental computational tool underpinning this analysis is the power flow problem, which is solved to obtain node voltages and line flows for a given set of nodal power injections [2]. Direct current power flow (DCPF) approximation is a popular approach as it provides a fast, unique, and reliable solution to the power flow problem [3, 4, 5, 6, 7, 8]. From a computational standpoint, the DCPF problem translates to solving a linear system of equations given a nodal power injection vector. This implies that with an increase in system size and changes in the injection vector due to uncertainty, the computational effort required to conduct a DCPF analysis increases. In addition to DCPF, another prominent linear power flow formulation called fast decoupled load flow (FDLF) also reduces to solving a sequence of linear systems of equations [2]. In practice, algorithms such as the conjugate gradient (CG) method are used to solve these linear power flow problems for large-scale systems, exploiting features such as sparsity for faster convergence to the solution [9].
@@
Quantum computing (QC), in theory, can provide provably asymptotic speedups over classical computers for many problems of practical interest [10, 11, 12]. QC achieves this by exploiting inherently quantum properties such as superposition and entanglement. This leads to novel algorithmic techniques in QC that cannot be emulated efficiently by classical computers. One such algorithmic primitive developed in QC literature is an exponentially fast linear system solver, first introduced in the work of Harrow, Hassidim, and Lloyd (HHL) [13] and improved upon by others [14, 15]. The exponential speed up in these works comes from the fact that linear system solving subroutine can solve a sparse linear system in time that scales only logarithmically with system size, if the input to the subroutine is provided as a quantum state and if relevant output information can be extracted using a small number of quantum measurements. The input/output model used in these algorithms is markedly different from those in classical computing. This indicates that a true end-to-end exponential speed-up cannot be claimed over classical algorithms in general. This point has been noted before in QC literature [16] and in some cases the claimed exponential speedup has been shown to vanish when the classical computer is given access to a data input/output model that appropriately matches the data model used by the QC algorithm [17, 18, 19].
@@
Figure 1: Power network data, equivalent linear system of equations, and the idea of practical quantum advantage (PQA). Left: Consider a simplified power grid with four “buses” (nodes) connected by five “branches” (lines). Each line has a certain “susceptance” which affects how easily electricity flows through it. Using direct current approximations [3], we can approximate the network by focusing only on these susceptances and ignoring other electrical properties (In the DC power flow approximation, the network is assumed to be lossless, the phase-angle differences between nodes are considered small, and all node voltages are assumed to be approximately equal to one per unit [3]). The unknowns in this system are the bus angle values (depicted in red) which depend on known quantities like susceptance (depicted in blue). Without any loss of generality, we can set one of the bus angle values to zero. Center: To solve the problem, DCPF converts the grid into a set of equations. It uses the susceptance values, known bus angles, and power injections (how much electricity enters or leaves each bus) to calculate the unknown bus angle values. It can be shown that the $A$ in this system is always positive by diagonal dominance arguments[3]. Right: Complexity analysis of classical and quantum approaches to this problem tells us how their respective runtimes scale with problem size and other parameters. PQA can exist only if there is a crossover point as depicted in this figure. Such a crossover implies that there exists a system size beyond which quantum algorithms can be faster for solving this problem.
@@
Figure 2: A simple information flow schematic of solving any problem using classical and quantum algorithms. From Figure 1, classical input is $\mathbf{b}$, and classical output is $\mathbf{x}$. The quantum algorithm encodes the matrix $A$ (considering $A$ as a Hamiltonian for a quantum system) in terms of time evolution operator $e^{iAt}$. As depicted here, solving a linear system using quantum computing via the HHL algorithm involves three steps [20]: i) state preparation; ii) state propagation (quantum algorithm), and iii) readout. In the state preparation step problem parameter $\mathbf{b}$ is encoded into a quantum state $|\Psi_{o}\rangle$. Following this, the second step is the application of quantum gates (unitary operations) to evolve the initial quantum state $|\Psi_{o}\rangle$ into final state $|\tilde{\textbf{x}}\rangle$ based on the properties of the matrix $A$. Here, $|\tilde{\textbf{x}}\rangle$ is a quantum state and the third step is readout, where $|\tilde{\textbf{x}}\rangle$ is measured to extract the solution to the linear system $\widehat{\mathbf{x}}$. We assess the quantum advantage based on the end-to-end run-time complexities for the DCPF problem. Here, end-to-end complexity in QC refers to the comprehensive time complexity from input to output, encompassing all stages of solving the DCPF problem, and not just the cost of running a quantum circuit that is a subroutine of the full solution method The end-to-end complexity determines the existence of PQA for these problems as it tells us how the total runtime required to solve this problem scales with the size of the system and other parameters in realistic DCPF problems.
@@
In recent times, QC has garnered significant attention in the power systems community, with efforts directed towards solving power flow problems through QC techniques, referred to as quantum power flow (QPF) [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. The primary rationale behind these endeavors lies in QC’s theoretical ability to efficiently solve a linear systems of equations. This is particularly pertinent as various power flow formulations, such as DCPF and FDLF, are either equivalent or can be reduced to solving linear systems of equations [3, 2]. In [21, 23, 22], the authors suggest employing the HHL algorithm to solve FDLF on a quantum computer, implying a potential exponential speed-up due to the runtime complexity of HHL. Authors in [23] present results obtained using a real quantum computer for solving the FDLF problem. The authors in [21] and [22] propose exponential speedup but the complexities associated with the readout of FDLF iteration outputs have not been addressed. This aspect, as highlighted in [23], is crucial for the execution of FDLF. Additionally, in [23], the authors rightly argue that quantum memory is indispensable for attaining quantum speedup in FDLF and suggest probabilistic Monte-Carlo simulation-type problems as potential applications of QPF. However, the paper does not analyze end-to-end complexity including the stages like state preparation. To alleviate qubit requirements, [28] proposes a hybrid method. Additionally, [30, 24] describes variational QPF to reduce the circuit depth needed to solve QPF compared to HHL-based QPF [21]. Authors in [24] provide a comprehensive review of different QPF formulations and highlight the need to multiply the ‘downloading and uploading’ complexities in HHL-QPF complexity, however, their effect on quantum speedup remains to be explored.
@@
Due to these limitations of existing discussions on QPF formulations there is a lack of a thorough assessment of the requirements for a practical quantum advantage (PQA). Quantum advantage refers to the potential computational superiority of quantum computers over classical computers for specific problems [32]. The challenge in the PQA assessment can be linked to plotting a runtime complexity or resource requirement curve, as illustrated in the rightmost subplot of Figure 1. In the context of DCPF, identifying the crossover point on this plot, that is, where the quantum computer’s resource requirements become more favorable than the best-known classical method, holds significant implications for the practical adoption of QC in power systems. If such a crossover point does not exist or if the crossover point exceeds the size of real-world problems (e.g., a billion-node power flow problem), the adoption of QC may not be practically justified. Specifically for the QPF problem, achieving PQA involves demonstrating that a quantum algorithm can efficiently provide the complete network state vector (including both the voltage angle and magnitude, depending on the formulation) by solving a power flow problem, significantly faster or with fewer computational resources and with the desired level of accuracy, than the best-known classical algorithms.
@@
In this paper, we elaborate on the complexities associated with each step of solving the QPF, shown in Figure 2. Further, the exponential speed-up claim of HHL-QPF methods is critically weighted against various HHL caveats [16]. Further, by doing the end-to-end complexity assessment for this problem we attempt to build a framework to assess:
@@
This work does not solve QPF, but instead hones the runtime complexity of QPF by considering the end-to-end time requirements and incorporating the key parameters from realistic AC power network datasets [35]. The main contributions of this paper can be summarized as:
  • Demonstration that the end-to-end runtime complexity of solving the DCPF problem through the HHL-QPF algorithm is asymptotically worse than that of the classical CG method. When considering the read-out [34] requirement in QC and condition number growth with network size ($N$ Buses) for realistic AC power network datasets [35], the complexity of HHL-QPF becomes $\mathcal{O}(N^{4.62}\log N)$, whereas the CG method’s complexity for solving an $N$-Bus DCPF problem remains $\mathcal{O}(N^{1.90}\log N)$. The effect of runtime overhead of state preparation is also discussed.
  • Identifying the power network parameter range (condition number) and readout level [^1] Readout level refers to the amount of classical information needed to extract the desired result from a quantum computation. For example, if we only need the mean of all states, the readout level is one, as we only require a single classical value. Consider a power system with N=100 buses. If we’re interested in the angles at buses 1, 20, 33, 55, and 99, in a 100-Bus system, the readout level would be five, even though the full vector $\mathbf{x}$ contains angles for 99 buses. for the types of the power flow-type problems for which practical quantum advantage can potentially exist. We show that for readout requirement of $D$ states and condition number $\kappa$, problem parameters must satisfy222 Where, $s$ is the sparsity of matrix $A$ defined as largest number of non-zero elements in a row. $D\kappa^{3/2}\ll Ns^{-1}$ to have any potential PQA using the HHL algorithm.
@@ II DC Power Flow Problem
The DCPF equations can be obtained by applying several simplifying assumptions to the non-linear alternating current (AC) branch flow equations [3]. These assumptions lead to conditions where the power balance equations can be expressed as a weighted sum of angle differences between buses. The weights in this sum are represented by the elements of the susceptance matrix as shown in Figure 1. In matrix form the DCPF is,
@@
with $A$ being positive definite, symmetric matrix of size $N\times N$ where $N$ is one less than system size, $\mathbf{x}$ and $\mathbf{b}$ are $N$-dimensional vectors. [^3] We have opted to use the standard linear system notation $A\mathbf{x}=\mathbf{b}$ instead of the standard DCPF notation $B\boldsymbol{\theta}=\mathbf{p}$. Here, $B$ is the susceptance matrix, $\boldsymbol{\theta}$ is the node phase-angle vector, and $\mathbf{p}$ is the nodal real-power injection vector. This choice facilitates compatibility with the analysis of linear system solving complexity for both classical and quantum paradigms.
@@
The runtime complexity to solve (1) using Gaussian elimination is $\mathcal{O}(N^{3})$ or precisely $\mathcal{O}(N^{\omega})$ where $\omega\in[2,2.372)$ denotes the matrix multiplication exponent [12]. This complexity becomes prohibitive in solving DCPF for large-scale systems because an industry scale network of 10,000 buses will require $\sim 10^{10}$ operations, which is daunting for classical computers processing operations at gigahertz speeds. Further, exact methods are not always able to exploit the sparsity present in system matrix $A$. In a power network, each node is connected to only a few other nodes of the network grid. Thus, sparsity is present in $A$ for all practical power networks. Figure 3 (left) shows the sparsity level for different systems from PGLib[35] library, which includes over sixty open-access transmission network datasets.
Further, a computational gain can come from the fact that we do not require the exact solution of $\mathbf{x}$ in DCPF problem. This is because the least count of the measurement devices in a power network is much higher than computing machine precision, and thus an error of the order $10^{-6}$ is generally sufficient.
@@
Figure 3: Network properties of the PGLib [35] power transmission network datasets reveal structural insights and potential numerical challenges in solving DCPF. Left: Level of sparsity ($s$) for different power system size ($N$-Buses). Here, average sparsity is the average number of non-zero elements in any row of matrix $A$. The PEGASE grids tend to have higher $s$ values due to sub-system equivalencies. Sparsity levels are below 30 for most of the 66 power grids from PGLib, with an average sparsity under 10, highlighting the potential for efficient sparse matrix techniques. Center: Polynomial relationship of condition number ($\kappa$) of $A$ with power system size ($N$ Buses). The one-norm is used to lower bound $\kappa$ for the 78484-Bus system due to computational restrictions of calculating the two-norm of such a large system. Condition number exhibits a polynomial (near-quadratic) growth trend with increasing system size ($N$-Buses), signaling potential numerical challenges and runtime increase in solving DCPF. Right: Percentage of non-zero elements in load vectors (sparsity percentage) for different PGLib systems. It shows that load vector sparsity will not be able to reduce the encoding complexity of $\mathbf{b}$ vector beyond a certain level. More importantly, due to the nature of power grid operation applications, one cannot make strong assumptions about or requirements on load-vector sparsity levels.

### v3

# Limitations of Fault-Tolerant Quantum Linear System Solvers for Quantum Power Flow

@@ Abstract
Quantum computers hold promise for solving problems intractable for classical computers, especially those with high time or space complexity. Practical quantum advantage can be said to exist for such problems when the end-to-end time for solving such a problem using a classical algorithm exceeds that required by a quantum algorithm. Reducing the power flow (PF) problem into a linear system of equations allows for the formulation of quantum PF (QPF) algorithms, which are based on solving methods for quantum linear systems such as the Harrow-Hassidim-Lloyd (HHL) algorithm. Speedup from using QPF algorithms is often claimed to be exponential when compared to classical PF solved by state-of-the-art algorithms. We investigate the potential for practical quantum advantage in solving QPF compared to classical methods on gate-based quantum computers. Notably, this paper does not present a new QPF solving algorithm but scrutinizes the end-to-end complexity of the QPF approach, providing a nuanced evaluation of the purported quantum speedup in this problem. Our analysis establishes a best-case bound for the HHL-based quantum power flow complexity, conclusively demonstrating that the HHL-based method has higher runtime complexity compared to the classical algorithm for solving the direct current power flow (DCPF) and fast decoupled load flow (FDLF) problem. Notably, our analysis and conclusions can be extended to any quantum linear system solver with rigorous performance guarantees, based on the known complexity lower bounds for this problem. Additionally, we establish that for potential practical quantum advantage (PQA) to exist it is necessary to consider DCPF-type problems with a very narrow range of condition number values and readout requirements.
@@ I Introduction
Providing electrical energy to the world and fighting climate change by switching from fossil fuel–powered to electric-powered transportation will require electricity generation to increase by more than two and half times by 2050 from 2021 levels [1]. This increased reliance on electricity will make secure power grid operation an even more important challenge. Modern power system operation relies heavily on the ability of system operators to accurately and efficiently compute the system state. The fundamental computational tool underpinning this analysis is the power flow problem, which is solved to obtain node voltages and line flows for a given set of nodal power injections [2]. Direct current power flow (DCPF) approximation is a popular approach as it provides a fast, unique, and reliable solution to the power flow problem [3, 4, 5, 6, 7, 8]. From a computational standpoint, the DCPF problem translates to solving a linear system of equations given a nodal power injection vector. Therefore, an increase in system size and changes to the injection vector due to uncertainty increases the computational effort required to conduct DCPF analysis. Another prominent linear power flow formulation, fast decoupled load flow (FDLF), also reduces to solving a sequence of linear systems of equations [2]. Solving fast DCPF and FDLF is imperative for secure power system operations under uncertainty, as they serve as essential subroutines in various probabilistic and stochastic analyses, such as probabilistic power flows. For instance, evaluating the percentage of lines violating power flow constraints over one million possible load points leads to a total runtime of $10^{6}\times\text{Single-instance Runtime}$. In such cases, even a marginal speedup in the runtime of a single instance of DCPF or FDLF can have a significant impact on the overall computational efficiency, especially when considering the large number of load samples involved, even with considerable parallelization. In practice, algorithms such as the conjugate gradient (CG) method are used to solve these linear power flow problems for large-scale systems, exploiting features such as sparsity for faster convergence to the solution [9].
@@
Quantum computing (QC), in theory, can provide provably asymptotic speedups over classical computers for many problems of practical interest [10, 11, 12]. QC achieves this by exploiting inherently quantum properties such as superposition and entanglement. This leads to novel algorithmic techniques in QC that cannot be emulated efficiently by classical computers. One such algorithmic primitive developed in QC literature is an exponentially fast linear system solver, first introduced in the work of Harrow, Hassidim, and Lloyd (HHL) [13] and improved on by others [14, 15]. Exponential speedup in these techniques comes from the fact that a linear system solving subroutine can solve a sparse linear system in a time that scales only logarithmically with system size, if subroutine input is provided as a quantum state and if relevant output information can be extracted using a small number of quantum measurements. The input/output model used in these algorithms is markedly different from those used in classical computing. This suggests that, in general, a true end-to-end exponential speedup cannot be claimed over classical algorithms. This point has been noted in QC literature [16], and in some cases the claimed exponential speedup has been shown to vanish when the classical computer is given access to a data input/output model that appropriately matches the data model used by the QC algorithm [17, 18, 19].
@@
Figure 1: Power network data, equivalent linear system of equations, and the idea of practical quantum advantage (PQA). Left: Consider a simplified power grid with four “buses” (nodes) connected by five “branches” (lines). Each line has a certain “susceptance” that affects how easily electricity flows through it. Using direct current approximations [3], we can approximate the network by focusing only on the susceptances and ignoring other electrical properties (in the DC power flow approximation, the network is assumed to be lossless, the phase-angle differences between nodes are considered small, and all node voltages are assumed to be approximately equal to one per unit [3]). The unknowns in this system are the bus angle values (shown in red), which depend on known quantities like susceptance (shown in black). Without any loss of generality, we can set one of the bus angle values to zero. Center: To solve the problem, DCPF converts the grid into a set of equations. It uses the susceptance values, known bus angles, and power injections (how much electricity enters or leaves each bus) to calculate the unknown bus angle values. Diagonal dominance arguments show that AA in this system is always positive definite [3, 20], and one can always construct a $A^{\dagger}A\mathbf{x}=A^{\dagger}\mathbf{b}$, where $A^{\dagger}$ is complex conjugate of matrix AA, and $A^{\dagger}A$ is Hermitian. Right: Complexity analysis of classical and quantum approaches to this problem tells us how the respective runtimes of these approaches scale with problem size and other parameters. PQA can exist only if there is a crossover point as depicted in this figure. Such a crossover implies that there exists a system size beyond which quantum algorithms can be faster for solving this problem.
@@
Figure 2: A simple information flow schematic of solving a problem using classical and quantum algorithms. From Figure 1, classical input is $\mathbf{b}$, and classical output is $\mathbf{x}$. The quantum algorithm encodes the matrix AA (considering AA as a Hamiltonian for a quantum system) in terms of the time evolution operator $e^{iAt}$. As depicted, solving a linear system using quantum computing by means of the HHL algorithm involves three steps [21]: (i) state preparation; (ii) state propagation (quantum algorithm), and (iii) readout. In the state preparation step problem, parameter $\mathbf{b}$ is encoded into a quantum state $|\Psi_{o}\rangle$. Following this, the second step is the application of quantum gates (unitary operations) to evolve the initial quantum state $|\Psi_{o}\rangle$ into the final state $|\tilde{\textbf{x}}\rangle$ based on the properties of the matrix $A$. Here, $|\tilde{\textbf{x}}\rangle$ is a quantum state and the third step is a readout, where $|\tilde{\textbf{x}}\rangle$ is measured to extract the solution to the linear system $\widehat{\mathbf{x}}$. We assess the quantum advantage based on the end-to-end runtime complexities of the DCPF problem. Here, end-to-end complexity in QC refers to the comprehensive time complexity from input to output, encompassing all stages of solving the DCPF problem and not just the cost of running a quantum circuit that is a subroutine of the full solution method. The end-to-end complexity determines the existence of PQA for these problems by showing how the total runtime required to solve a problem scales with system size and other parameters in realistic DCPF problems.
@@
Recently, QC has garnered significant attention in the power systems community, with efforts directed towards solving power flow problems through QC techniques, referred to as quantum power flow (QPF) [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. The primary rationale behind these endeavors lies in QC’s theoretical ability to efficiently solve a linear systems of equations. This is particularly pertinent as various power flow formulations, such as DCPF and FDLF, are either equivalent to or can be reduced to solving linear systems of equations [3, 2]. In [22, 24, 23], the authors suggest employing the HHL algorithm to solve FDLF on a quantum computer, implying a potential exponential speedup due to the runtime complexity of HHL. Authors in [24] present results obtained using a real quantum computer for solving the FDLF problem. The authors in [22] and [23] propose exponential speedup, but the complexities associated with the readout of FDLF iteration outputs are not addressed. This aspect, as highlighted in [24], is crucial for the execution of FDLF. Additionally, in [24], the authors rightly argue that quantum memory is indispensable for attaining quantum speedup in FDLF and suggest probabilistic Monte-Carlo simulation-type problems as potential applications of QPF. However, the paper does not analyze end-to-end complexity, which includes stages like state preparation. To alleviate qubit requirements, [29] proposes a hybrid method. Additionally, [31, 25] describe variational QPF to reduce the circuit depth needed to solve QPF compared to that needed for HHL-based QPF [22]. Authors in [25] provide a comprehensive review of different QPF formulations and highlight the need to multiply the ‘downloading and uploading’ complexities in HHL-QPF complexity. However, the effect of these complexities on quantum speedup remains to be explored.
@@
Overall a through assessment of the requirements for quantum advantage is lacking in the current literature. Quantum advantage refers to the potential computational superiority of quantum computers over classical computers for specific problems [33]. The challenge of a PQA assessment can be shown by plotting a runtime complexity or resource requirement curve, as illustrated in the rightmost subplot of Figure 1. In the context of DCPF, identifying the crossover point on this plot, that is, where the quantum computer’s resource requirements become more favorable than the best-known classical method, holds significant implications for the practical adoption of QC in power systems. If such a crossover point does not exist or if the crossover point exceeds the size of real-world problems (e.g., a billion-node power flow problem), the adoption of QC may not be practically justified. For the QPF problem in specific, achieving PQA involves demonstrating that a quantum algorithm can efficiently provide the complete network state vector (including both the voltage angle and magnitude, depending on the formulation) by solving a power flow problem significantly faster or with fewer computational resources than and with the desired level of accuracy of the best-known classical algorithms.
@@
In this paper, we elaborate on the complexities associated with each step of solving the QPF, shown in Figure 2. Further, we critically weigh the exponential speedup claim of HHL-QPF methods against various HHL caveats [16]. Further, by completing an end-to-end complexity assessment for this problem, we attempt to build a framework to assess the following:
@@
It is important to highlight that this work does not provide a new QPF solving algorithm, but instead hones in on the runtime complexity of QPF by considering end-to-end time requirements and incorporating key parameters from realistic AC power network datasets [34]. The main contributions of this paper can be summarized as follows:
  • Demonstrating that the end-to-end runtime complexity of solving the DCPF problem using the HHL-QPF algorithm is asymptotically worse than that of the classical CG method. When considering the readout [35] requirement in QC and condition number growth with network size ($N$ Buses) for realistic AC power network datasets [34], the complexity of HHL-QPF becomes $\mathcal{O}(N^{4.62}\log N)$, whereas the CG method’s complexity for solving an $N$-Bus DCPF problem remains $\mathcal{O}(N^{1.90}\log N)$. The effect of runtime overhead of state preparation is also discussed.
  • Identifying the power network parameter range (condition number) and readout level [^1] Readout level refers to the amount of classical information needed to extract the desired result from a quantum computation. For example, if we only need the mean of all states, the readout level is one, as we only require a single classical value. Consider a power system with N=100 buses. If we’re interested in the angles of buses 1, 20, 33, 55, and 99 in a 100-bus system, the readout level would be five, even though the full vector $\mathbf{x}$ contains angles for 99 buses. for the power flow–type problems for which PQA can potentially exist. We show that for readout requirements of $D$ states with condition number $\kappa$, problem parameters must satisfy [^2] Where, condition number $\kappa$ is the ratio of largest to smallest eigenvalue and $s$ is the sparsity of matrix $A$ defined as largest number of non-zero elements in a row. $D\kappa^{3/2}\ll Ns^{-1}$ to have any potential PQA using the HHL algorithm.
Remark 1.
Unlike HHL and its improvements, the Variational Quantum Linear Solver (VQLS) is a hybrid quantum-classical algorithm [36], where ansatz is updated using a hybrid quantum-classical optimization loop, to achieve local or global cost below user-specified threshold. This makes VQLS a heuristic algorithm, unlike HHL, which is a purely quantum algorithm with the stopping criterion defined by the measurement of an ancilla bit (Appendix B. Brief Explanation of HHL discusses this criteria of algorithm success) [13, 37]. Therefore, the end-to-end runtime complexity discussion does not apply to VQLS. Furthermore, due to its heuristic nature, an empirical study is required to evaluate the effectiveness of VQLS in solving DCPF (DC Power Flow) problems and to understand its practical applicability. It is worth noting that the paper introducing VQLS [36] only conducts such a study to provide an upper bound on the quality of the solution, not on the runtime.
Remark 2.
We only consider quantum linear system solvers designed to run on a error corrected, fault tolerant quantum computers. Most theoretical claims of quantum advantage are made for this type of quantum computers. While quantum algorithms can be run on Noisy Intermediate Scale Quantum (NISQ) devices, these do not come with rigorous performance guarantees and can be limited in terms of scalability [38, 39].
@@ II DC Power Flow Problem
The DCPF equations can be obtained by applying several simplifying assumptions to the non-linear alternating current (AC) branch flow equations [3]. These assumptions lead to conditions in which the power balance equations can be expressed as a weighted sum of the angle differences between buses. The weights in this sum are represented by the elements of the susceptance matrix as shown in Figure 1. In matrix form, the DCPF is
@@
with $A$ being positive definite, symmetric matrix of size $N\times N$ where $N$ is one less than system size, and $\mathbf{x}$ and $\mathbf{b}$ are $N$-dimensional vectors. [^3] We have opted to use the standard linear system notation $A\mathbf{x}=\mathbf{b}$ instead of the standard DCPF notation $B\boldsymbol{\theta}=\mathbf{p}$. Here, $B$ is the susceptance matrix, $\boldsymbol{\theta}$ is the node phase-angle vector, and $\mathbf{p}$ is the nodal real-power injection vector. This choice facilitates compatibility with the analysis of linear system solving complexity for both classical and quantum paradigms.
@@
The runtime complexity to solve (1) using Gaussian elimination is $\mathcal{O}(N^{3})$, or, precisely, $\mathcal{O}(N^{\omega})$, where $\omega\in[2,2.372)$ denotes the matrix multiplication exponent [12]. This complexity becomes prohibitive when solving DCPF for large-scale systems because an industry scale network of 10,000 buses will require $\sim 10^{10}$ operations, which is daunting for classical computers processing operations at gigahertz speeds. Further, exact methods matrix inversion (e.g. Gaussian Elimination) are not always able to exploit the sparsity present in system matrix $A$. In a power network, each node is connected to only a few other nodes of the network grid. Thus, sparsity is present in $A$ for all practical power networks. Figure 3 (left) shows the sparsity level for different systems from PGLib[34] library, which includes over sixty open-access transmission network datasets.
Moreover, the computational efficiency can be improved by not necessitating the precise solution of $\mathbf{x}$ in the DCPF problem. Typically, errors within the range of $10^{-4}$ to $10^{-6}$ are considered reasonable, varying based on the specific application requirements. This range is deemed acceptable due to the resolution limitations of physical measurement devices within a power network. The impact of the error value on runtime is quantified by the parameter $\varepsilon$ in our complexity discussions.
@@
Figure 3: Network properties of the PGLib [34] power transmission network datasets reveal structural insights and potential numerical challenges in solving DCPF. Left: Level of sparsity ($s$) for different power system size (NN-buses). Here, average sparsity is the average number of non-zero elements in any row of matrix $A$. The PEGASE grids tend to have higher $s$ values due to subsystem equivalencies. Sparsity levels are below 30 for most of the $6$ power grids from PGLib, with an average sparsity under 10, highlighting the potential for efficient sparse matrix techniques. Center: Polynomial relationship of condition number ($\kappa$) of $A$ with power system size ($N$ buses). The one-norm is used to lower bound $\kappa$ for the 78484-Bus system due to computational restrictions of calculating the two-norm of such a large system. Condition number exhibits a polynomial (near-quadratic) growth trend with increasing system size (NN-buses), signaling potential numerical challenges and runtime increase in solving DCPF. Right: Percentage of non-zero elements in injection vectors (sparsity percentage) for different PGLib systems. It shows that injection vector ($\mathbf{b}$ in DCPF formulation) sparsity will not be able to reduce the encoding complexity of $\mathbf{b}$ vector beyond a certain level. More importantly, due to the nature of power grid operation applications, one cannot make strong assumptions about or requirements on load-vector sparsity levels. We use PowerModels.jl[40] to construct the susceptance matrix for all these PGLib [34] systems.